OpenStudy (anonymous):

Perform the operations & write result in standard form. (4+3i)(2-5i) Is this done by distribution?

OpenStudy (anonymous):

use foil

OpenStudy (anonymous):

thought so.

OpenStudy (anonymous):

but watch-- the multipication of i can be a little tricky

OpenStudy (anonymous):

how so. Isn't it a matter on combining like terms?

OpenStudy (anonymous):

Have so far 8-20i=6i-15i=

OpenStudy (anonymous):

should be 8-20i+6i-15i

OpenStudy (anonymous):

yes but what do you get here 8-20i+6i-15i^2 8-14i+15 23-14i

OpenStudy (anonymous):

you forgot to square the i and i^2=-1

OpenStudy (anonymous):

So.... answer should be 8-14i+15

OpenStudy (anonymous):

yes but standar form is a-bi so8+15=23 so 23-14i

OpenStudy (anonymous):

ok